Digital Healthcare Co., a mid-size electronic medical device maker, has decided to downsize as a measure to reduce cost in response weak demand. The company is considering implementing one of two plans:

the first plan involves closing and selling a manufacturing plant in Maple Ridge, which is expected to generate a salvage value of $5 million today (year 0) and save $1.5 million annually in cost and expenses over the next 10 years (years 1 to 10). However, the company will lose $1 million in revenue the first year (year 1), and the loss will increase by 10% per year over the following 9 years (years 2 to 10).

NPV Plan A = $____ million
Answer #1

Net Present Value = PV of cash inflows- PV of cash outflows

Assuming Discounting Rate is 10%

Salvage value that can be earned Now (Year0)= $5 Million(Cash inflow) (A)

As $1.5 million is saved every year ,it is similar to cash inflow and Present cash flows of $1.5 million is to be calculated for 10 years @10% rate.

Discounting factor is calculated as follows: 1/(1+R)^n

Where R is Discounting rate, N is number of years for which it has to be discounted.

Discounting factors @10% rate for 10 years are as follows

Year 1 2 3 4 5 6 7 8 9 10
Discounting factor 0.909 0.826 0.751 0.683 0.621 0.564 0.513 0.467 0.424 0.386

  

Sum of discounting factors from 1 to 10 years is (0.909+0.826+....+0.386) =6.145

Therefore present value of $1.5 million saved annualy for 10 years is = $1.5 million * 6.145

=$9216850.65 (B)

Losses to be incurred by company for 10 years are as follows: $ in millions

Year 1 2 3 4 5 6 7 8 9 10
$ in milions 1 1.1 1.2 1.3 1.4 1.5 1.6 1.7 1.8 1.9

As loss increases every year by 10% , Loss is varying every year , this is similar to cash outflow.

Now, Calculate present value of above losses for 10 years

Year 1 2 3 4 5 6 7 8 9 10
Loss in $millions (a) 1 1.1 1.2 1.3 1.4 1.5 1.6 1.7 1.8 1.9
Disc. Factor@ 10% (b) 0.909 0.826 0.751 0.683 0.621 0.564 0.513 0.467 0.424 0.386
Present Value (a*b) 0.909 0.909 0.902 0.888 0.869 0.847 0.821 0.793 0.763 0.733

Sum of above present values (0.909+0.909+....+0.733)= $8.434 millions (C)

Therefore net present value is = (A) +(B) -(C)

=5000000+9216850.65-8434000

=$5783149
Answer #2

                 

  • Plan A:

    NPV=5+1.50.12(111.1210)10.120.1(1(1.11.12)10)=$5.23 million

  • Plan B:

    NPV=811.12530.12(111.125)=$1.86 million

Plan A has a larger NPV.

         


answered by: palak